# SARCLIP: A Vision Language Foundation Model for Semantic Understanding and Target Recognition in SAR Imagery

### v1

@@ Abstract
Synthetic Aperture Radar (SAR) has emerged as a crucial imaging modality due to its all-weather capabilities. While recent advancements in self-supervised learning and Masked Image Modeling (MIM) have paved the way for SAR foundation models, these approaches primarily focus on low-level visual features, often overlooking multimodal alignment and zero-shot target recognition within SAR imagery. To address this limitation, we construct SARCLIP-1M, a large-scale vision language dataset comprising over one million text-image pairs aggregated from existing datasets. We further introduce SARCLIP, the first vision language foundation model tailored for the SAR domain. Our SARCLIP model is trained using a contrastive vision language learning approach by domain transferring strategy, enabling it to bridge the gap between SAR imagery and textual descriptions. Extensive experiments on image-text retrieval and zero-shot classification tasks demonstrate the superior performance of SARCLIP in feature extraction and interpretation, significantly outperforming state-of-the-art foundation models and advancing the semantic understanding of SAR imagery. The code and datasets will be released soon.
I Introduction
SAR is a high-resolution imaging technology with all-weather, day-night operability and strong penetration capabilities, widely applied in military, environmental, maritime, and disaster monitoring tasks. Compared with optical imagery, SAR images are characterized by speckle noise, geometric distortions, and limited semantic textures, which pose significant challenges for downstream tasks such as ship detection [1], aircraft recognition [2], object detection [3], semantic segmentation [4], and image classification [5]. These challenges highlight the need for robust and generalizable feature representations specifically tailored to SAR data.
Figure 1: Comparison of image-text datasets in remote sensing. Gray circles represent datasets in the optical image domain, blue circles denote those in the SAR domain, and the red circle indicates our proposed SARCLIP-1M. In the SAR domain, SARCLIP-1M significantly surpasses existing datasets in both image and text volume.
Recent developments in vision foundation models (VFMs) have led to promising generalization capabilities across domains. As illustrated in Fig. 2, existing VFMs can be broadly categorized into three paradigms: (a) contrastive learning (CL-based) methods such as SimCLR [6], which aim to learn discriminative representations by aligning augmented views of the same image; (b) masked image modeling (MIM-based) methods that reconstruct occluded image regions to learn spatial representations; and (c) CLIP-based approaches [7] that align images and texts via contrastive loss on large-scale paired data, showing remarkable transferability in cross-modal tasks.
@@ I Introduction
The self-supervised training paradigm has recently driven the development of several vision foundation models in remote sensing, including RS-BYOL [8], ScaleMAE [9], and Cross-Scale MAE [10]. Notably, ScaleMAE [9] focuses on learning multi-scale representations by reconstructing both low- and high-frequency components across defined spatial scales. Similarly, Cross-Scale MAE [10] enforces consistency across scales using a combination of contrastive and generative losses to enhance self-supervised remote sensing representations. However, these methods predominantly address the extraction of visual features. Inspired by CLIP [7], RemoteCLIP [11] pioneered the first vision language foundation model (VLFM) in remote sensing, aiming to capture semantic features by aligning image-text pairs. Subsequent works like RS5M [12] and SkyScript [13] have further explored this aspect. However, these models are pre-trained on optical image, leaving the SAR modality underexplored due to its distinct characteristics and lack of large-scale image-text datasets in SAR domain.
In recent years, the growing interest in SAR modality has led to the creation of various large-scale SAR datasets [14, 15, 16, 17]. Specifically, SARDet-100K consolidates existing SAR datasets to form a large collection encompassing ships, aircraft, cars, bridges, tanks, and harbors. Building upon this, SARATR-X extends SARDet-100K by integrating additional classification datasets to create the larger SARDet-180K. Both SARDet-100K [14] and SARATR-X [15] then employ MIM training strategy to establish their respective foundation models. Furthermore, recent advanced methods have to address the application of multi-modal LLMs in the SAR field. Particularly, SARChat-2M [18] introduces a substantial benchmark with two million multimodal dialogues, facilitating intelligent interpretation of SAR imagery through LLM-based conversational paradigms. A comparable initiative has been undertaken with SARLANG-1M [19]. However, despite their impressive performance, existing SAR vision foundation models and multimodal LLMs remain limited by the lack of textual annotations in current datasets, hindering their ability to fully capture rich semantic information.
To bridge this gap, we propose SARCLIP, the first CLIP-based vision-language foundation model designed specifically for SAR imagery. As shown in Fig. 1, we construct SARCLIP-1M, a large-scale SAR image-text dataset comprising 1.7 million pairs across diverse object categories and land cover types. These pairs are generated by leveraging domain knowledge, spatial rules, and templated text synthesis strategies. Based on SARCLIP-1M dataset, we adopt a two-stage domain strategy to transfer knowledge from optical to SAR domain, enabling semantic alignment between modalities. As shown in Fig. 3, this approach enables the extraction of comprehensive general features from SAR images and significantly enhances their semantic understanding, thereby elevating performance for SAR interpretation.
@@
Figure 3: (a) Examples from the SARCLIP-1M dataset; (b) Domain transfer training strategy for SARCLIP.
The key contributions of our SARCLIP are summarized as follows:
@@
  • We propose SARCLIP-1M, a novel large-scale vision-language dataset containing 1.7 million image-text pairs, encompassing various object and land cover types.
@@
  • To the best of our knowledge, SARCLIP is the first VLFM tailored for SAR imagery, enabling SAR-specific cross-modal alignment, target recognition and zero-shot classification.
@@
  • Extensive experiments demonstrate that SARCLIP consistently surpasses existing state-of-the-art VLFMs on multiple downstream tasks, highlighting its strong generalization and semantic representation capabilities in the SAR domain.
II Related Work
II-A Vision Language Model for Remote Sensing
@@
Vision-language models have significantly advanced remote sensing through the development of multi-modal techniques. These methods broadly fall into two categories: contrastive methods and generative methods. Among contrastive approaches, CLIP [7] stands out as a pioneering work, employing a two-tower architecture to align visual and language features through extensive data from Internet. Inspired by CLIP, several studies have adapted this paradigm to remote sensing, yielding models such as RemoteCLIP [11], GeoRSCLIP [12], SkyScript [13], RSMCLIP [20] and Mall et al. [21]. Notably, RemoteCLIP [11] enhances pre-training by converting detection and segmentation annotations into image captions, facilitating CLIP-style contrastive learning for image-text alignment in remote sensing. In the realm of generative methods, works like GeoChat [22], EarthGPT [23], and LHRS-Bot [24] implement auto-regressive large language model (LLM) architectures for vision-text alignment. For instance, GeoChat [22] introduces a multimodal model with LLaVa [25] architecture in remote sensing, achieving multi-granularity alignment through parameter-efficient fine-tuning. However, most of these existing methods primarily focus on the optical modality in remote sensing, largely overlooking investigations in the SAR domain.
II-B SAR Foundation Model
@@
Foundation models, by pretraining on large-scale data, are capable of capturing generalizable visual representations that effectively support a wide array of downstream tasks. In remote sensing, recent advancements have seen the utilization of self-supervised learning and MIM techniques, as demonstrated by approaches such as RingMo [26], SatMAE [27], ScaleMAE [9], OmniSat [28], and Cross-Scale MAE [10]. These models have been widely applied in tasks like aerial object detection and target recognition.
@@
With the increasing availability of SAR imagery, a multitude of datasets have emerged, including SARDet-100K [14], SARATR-X [15], SAR-JEPA [29], FAIR-CSAR [16], and ATRNet-STAR [17]. Specifically, SARDet-100K constructs SAR foundation model through MIM training, initially pretraining on aerial view images before transferring to SAR imagery. Building upon this, SARATR-X leverages various classification datasets to build the SARDet-180K dataset, subsequently yielding a MIM-based SAR foundation model. Furthermore, SARLANG-1M [19] utilizes LLMs for SAR image interpretation, while SARChat [18] supports key tasks such as visual understanding and object detection in SAR imagery. Nevertheless, a common limitation among these existing SAR foundation models is their primary focus on low-level image features, often failing to capture deeper semantic information and realize multi-modal alignment within SAR images.
@@
TABLE I: Illustration of SARCLIP-1M, which includes 8 SAR datasets. Cls:Classification. Det.:Detection. Cap.:Caption. # Train Imgs.: Number of training imges. # Train Caps.: Number of training captions. # Val Imgs.: Number of validation images. # Val Caps.: Number of validation captions. # Test Pairs: Number of image-text pairs in testing set.
@@
MSTAR [30] 1995 Cls. 3,046 15,230 9,855 49,275 180
SARSim [31] 2017 Cls. 21,168 105,840
OpenSARShip [32] 2017 Cls. 26,679 133,395
SAMPLE [33] 2019 Cls. 5,380 26,900
ATRNet-STAR [17] 2025 Cls.&Det. 68,091 340,455 29,284 146,420 6,667
SARDet-100K [14] 2024 Det. 94,493 472,465 10,492 52,460 2,783
FAIR-CSAR [16] 2024 Det. 51,948 259,740 11,790 58,950 7,096
SARLANG-1M-Captions [19] 2025 Cap. 9,191 31,968 3,939 13,682 3,902
SARCLIP-1M(ours) 2025 Cap. 279,996 1,385,993 65,360 320,787 20,628
TABLE II: Overview of the datasets comprising SARCLIP-1M dataset. The three datasets1 are SpaceNet6 [34], DFC2023 [35], and OpenEarthMap [36]
. Datasets Descriptions MSTAR Contains X-band SAR imagery of military vehicles. SARSim A simulation dataset providing vehicle samples across 7 categories. OpenSARShip Features ship slices derived from European C-band Sentinel-1 satellite data. SAMPLE A public X-band SAR dataset of 10 vehicle classes, with synthetic and real image pairs. ATRNet-STAR A large-scale SAR dataset offering 40 fine-grained vehicle target classes. SARDet-100K Compiled from 10 existing SAR detection datasets, encompassing 5 object classes. FAIR-CSAR A large-scale, fine-grained SLC SAR dataset covering 22 subcategories. SARLANG-1M-Captions Contains over 45,000 SAR image captions based on three datasets1. SARCLIP-1M (ours) Comprises over 1.7 million image-text pairs, including ship, vehicle, aircraft, and other land covers.
III Methodology
@@
This section introduce the paradigm of our framework, dataset construction approach and training strategy for SARCLIP.
III-A Problem Definition
@@
In this section, we investigate the paradigm of learning joint representations from SAR images and their corresponding textual descriptions. Specifically, we construct SARCLIP-1M dataset $\mathcal{D}=\{(\mathbf{I}_{i},\mathbf{T}_{i})\}_{i=1}^{M}$ consisting of SAR images $\mathbf{I}_{i}\in\mathcal{R}^{H\times W}$ with the corresponding descriptions $\mathbf{T}_{i}\in\mathcal{T}$. As shown in Fig. 3, our objective is to learn a pair of modality-specific encoders that project both SAR images and text into a shared semantic space. Specifically, we define a visual encoder $f_{v}:\mathcal{R}^{H\times W}\rightarrow\mathcal{R}^{d}$ that maps the input SAR image to a $d$-dimensional visual feature embedding $z_{v}^{i}=f_{v}({\mathbf{I}_{i}})$, and a textual encoder $f_{t}:\mathcal{R}^{\mathcal{T}}\rightarrow\mathcal{R}^{d}$ that maps the textual input to a corresponding textual embedding $z_{t}^{i}=f_{t}({\mathbf{T}_{i}})$. The goal is to align the embeddings $z_{v}^{i}$ and $z_{t}^{i}$ of matched image-text pairs in a common representation space, such that semantically similar inputs across modalities are embedded close to each other. This formulation enables the model to bridge the modality gap between SAR images and natural language, thereby supporting a wide range of downstream tasks such as cross-modal retrieval, target recognition, and zero-shot classification in SAR domains.
III-B SARCLIP-1M Dataset Construction
@@
To address the aforementioned challenges, we build a large-scale vision-language dataset SARCLIP-1M collected from existing classification, detection and caption datasets in SAR domain. As illustrated in Table I and II, several large datasets have released in the past two year. For classification datasets, we employ 10 simple and complex description template to generate image captions. As for detection, we additionally design template to describe object absolute position and relative spatial position between targets in the images. The details for template design method are as follows:
@@
Based on the above templates design, we obtain caption for each SAR imagery. Then, A large language model is utilized to verify the fluency and grammatical correctness of text. Finally, we build our SARCLIP-1M dataset contains 279,996 images and 1,385,993 captions in training set, and 20,628 image-text pairs in testing set. As shown in Fig. 4, the dataset includes diverse target types such as ship, vehicle, aircraft, bridge and other land cover types.
III-C Model Training
III-C1 SARCLIP training.
To further enhance the semantic understanding ability of the model for SAR image, we utilize two stages training strategy to train our SARCLIP. Firstly, we train vision large model on natural remote sensing dataset like RemoteCLIP which contains LoveDA [37], DOTA [38], RSCID [39] and others. By this strategy, our model is capable of better understanding and recognizing the unique characteristics of remote sensing imagery, particularly in identifying small scale and densely distributed objects. Secondly, we finetune this pretrained model on our SARCLIP-1M dataset to transfer knowledge from natural to SAR domain.

### v2

# SARVLM: A Vision Language Foundation Model for Semantic Understanding and Target Recognition in SAR Imagery

@@ Abstract
Synthetic Aperture Radar (SAR) is a crucial imaging modality thanks to its all-weather capability. Although recent advances in self-supervised learning and masked image modeling (MIM) have enabled SAR foundation models, these methods largely emphasize low-level visual features and often overlook multimodal alignment and zero-shot target recognition in SAR imagery. To address this, we construct SARVLM-1M, a large-scale vision–language dataset with over one million image–text pairs aggregated from existing datasets. We further propose a domain transfer training strategy to mitigate the large gap between natural and SAR imagery. Building on this, we develop SARVLM, the first vision language foundation model (VLM) tailored to SAR, comprising SARCLIP and SARCap. SARVLM is trained with a vision–language contrastive objective under the proposed domain transfer strategy, bridging SAR imagery and textual descriptions. Extensive experiments on image text retrieval, zero-shot classification, semantic localization, and imagery captioning demonstrate that SARVLM delivers superior feature extraction and interpretation, outperforming state-of-the-art VLMs and advancing SAR semantic understanding. Code and datasets will be released soon.
1 Introduction
SAR is a high-resolution imaging technology with all-weather, day-night operability and strong penetration capabilities, widely applied in military, environmental, maritime, and disaster monitoring tasks. Compared with optical imagery, SAR images are characterized by speckle noise, geometric distortions, and limited semantic textures, which pose significant challenges for downstream tasks such as ship detection [SSDD], aircraft recognition [SFRNet], object detection [RSAR], semantic segmentation [liu2025learning], and image classification [deng2024hyperspectral]. These challenges highlight the need for robust and generalizable feature representations specifically tailored to SAR data.
Figure 1: Comparison of image-text datasets in remote sensing. Gray circles represent datasets in the optical image domain, blue circles denote those in the SAR domain, and the red circle indicates our proposed SARVLM-1M. In the SAR domain, SARVLM-1M significantly surpasses existing datasets in both image and text volume.
Recent developments in vision foundation models (VFMs) have led to promising generalization capabilities across domains. As illustrated in Figure 2, existing VFMs can be broadly categorized into three paradigms: (a) contrastive learning (CL-based) methods such as SimCLR [SimCLR], which aim to learn discriminative representations by aligning augmented views of the same image; (b) masked image modeling (MIM-based) methods that reconstruct occluded image regions to learn spatial representations; and (c) CLIP-based approaches [CLIP] that align images and texts via contrastive loss on large-scale paired data, showing remarkable transferability in cross-modal tasks.
@@ 1 Introduction
The self-supervised training paradigm has recently driven the development of several vision foundation models in remote sensing, including RS-BYOL [RS-BYOL], ScaleMAE [ScaleMAE], and Cross-scale MAE [Cross-Sclae-MAE]. Notably, ScaleMAE [ScaleMAE] focuses on learning multi-scale representations by reconstructing both low- and high-frequency components across defined spatial scales. Similarly, Cross-scale MAE [Cross-Sclae-MAE] enforces consistency across scales using a combination of contrastive and generative losses to enhance self-supervised remote sensing representations. However, these methods predominantly address the extraction of visual features. Inspired by CLIP [CLIP], RemoteCLIP [RemoteCLIP] pioneered the first vision language foundation model (VLFM) in remote sensing, aiming to capture semantic features by aligning image-text pairs. Subsequent works like RS5M [RS5M] and SkyScript [Skyscript] have further explored this aspect. However, these models are pre-trained on optical image, leaving the SAR modality underexplored due to its distinct characteristics and lack of large-scale image-text datasets in SAR domain.
@@
Figure 3: (a) Examples from the SARVLM-1M dataset; (b) Two-stage domain transfer training strategy for SARCLIP.
In recent years, the growing interest in SAR modality has led to the creation of various large-scale SAR datasets [SARDet-100K, SARATR, FAIR-CSAR, ATRNet-STAR]. Specifically, SARDet-100K consolidates existing SAR datasets to form a large collection encompassing ships, aircraft, cars, bridges, tanks, and harbors. Building upon this, SARATR-X extends SARDet-100K by integrating additional classification datasets to create the larger SARDet-180K. Both SARDet-100K [SARDet-100K] and SARATR-X [SARATR] then employ MIM training strategy to establish their respective foundation models. Furthermore, recent advanced methods have to address the application of multi-modal LLMs in the SAR field. Particularly, SARChat-2M [SARChat-Bench-2M] introduces a substantial benchmark with two million multimodal dialogues, facilitating intelligent interpretation of SAR imagery through LLM-based conversational paradigms. A comparable initiative has been undertaken with SARLANG-1M [SARLANG-1M]. However, despite their impressive performance, existing SAR vision foundation models and multimodal LLMs remain limited by the lack of textual annotations in current datasets, hindering their ability to capture rich semantic information fully.
To bridge this gap, we propose SARVLM consisting of SARCLIP and SARCap, the first vision-language foundation model designed specifically for SAR imagery. As shown in Figure 1, we construct SARVLM-1M, a large-scale SAR image-text dataset comprising 1.7 million pairs across diverse object categories and land cover types. These pairs are generated by leveraging domain knowledge, spatial rules, and templated text synthesis strategies. Based on SARVLM-1M dataset, we adopt a two-stage domain strategy to transfer knowledge from optical to SAR domain, enabling semantic alignment between modalities. As shown in Figure 3, this approach enables the extraction of comprehensive general features from SAR images and significantly enhances their semantic understanding, thereby elevating performance for SAR interpretation.
The key contributions of our SARVLM are summarized as follows:
@@
  • We propose SARVLM-1M, a novel large-scale vision-language dataset containing 1.7 million image-text pairs, encompassing various object and land cover types.
@@
  • To the best of our knowledge, SARVLM is the first VLFM tailored for SAR imagery, enabling SAR-specific cross-modal alignment, target recognition, zero-shot classification, semantic localization, and imagery captioning.
@@
  • Extensive experiments demonstrate that SARVLM consistently surpasses existing state-of-the-art VLFMs on multiple downstream tasks, highlighting its strong generalization and semantic representation capabilities in the SAR domain.
2 Related Work
2.1 Vision Language Model for Remote Sensing
@@
Vision-language models have significantly advanced remote sensing through the development of multi-modal techniques. These methods broadly fall into two categories: contrastive methods and generative methods. Among contrastive approaches, CLIP [CLIP] stands out as a pioneering work, employing a two-tower architecture to align visual and language features through extensive data from Internet. Inspired by CLIP, several studies have adapted this paradigm to remote sensing, yielding models such as RemoteCLIP [RemoteCLIP], GeoRSCLIP [RS5M], SkyScript [Skyscript], RSMCLIP [rethinking] and Mall et al. [mall2023remote]. Notably, RemoteCLIP [RemoteCLIP] enhances pre-training by converting detection and segmentation annotations into image captions, facilitating CLIP-style contrastive learning for image-text alignment in remote sensing. In the realm of generative methods, works like GeoChat [Geochat], EarthGPT [EarthGPT], and LHRS-Bot [LHRS-Bot] implement auto-regressive large language model (LLM) architectures for vision-text alignment. For instance, GeoChat [Geochat] introduces a multimodal model with LLaVa [LLaVa] architecture in remote sensing, achieving multi-granularity alignment through parameter-efficient fine-tuning. However, most of these existing methods primarily focus on the optical modality in remote sensing, largely overlooking investigations in the SAR domain.
2.2 SAR Foundation Model
@@
Foundation models, by pretraining on large-scale data, are capable of capturing generalizable visual representations that effectively support a wide array of downstream tasks. In remote sensing, recent advancements have seen the utilization of self-supervised learning and MIM techniques, as demonstrated by approaches such as RingMo [RingMo], SatMAE [SatMAE], ScaleMAE [ScaleMAE], OmniSat [OmniSat], and Cross-scale MAE [Cross-Sclae-MAE]. These models have been widely applied in tasks like aerial object detection and target recognition.
@@
With the increasing availability of SAR imagery, a multitude of datasets have emerged, including SARDet-100K [SARDet-100K], SARATR-X [SARATR], SAR-JEPA [SAR-JEPA], FAIR-CSAR [FAIR-CSAR], and ATRNet-STAR [ATRNet-STAR]. Specifically, SARDet-100K constructs SAR foundation model through MIM training, initially pretraining on aerial view images before transferring to SAR imagery. Building upon this, SARATR-X leverages various classification datasets to build the SARDet-180K dataset, subsequently yielding a MIM-based SAR foundation model. Furthermore, SARLANG-1M [SARLANG-1M] utilizes LLMs for SAR image interpretation, while SARChat [SARChat-Bench-2M] supports key tasks such as visual understanding and object detection in SAR imagery. Nevertheless, a common limitation among these existing SAR foundation models is their primary focus on low-level image features, often failing to capture deeper semantic information and realize multi-modal alignment within SAR images.
@@
Table 1: Illustration of SARVLM-1M, which includes 8 SAR datasets. Cls:Classification. Det.:Detection. Cap.:Caption. # Train Imgs.: Number of training imges. # Train Caps.: Number of training captions. # Val Imgs.: Number of validation images. # Val Caps.: Number of validation captions. # Test Pairs: Number of image-text pairs in testing set.
@@
MSTAR [MSTAR] 1995 Cls. 3,046 15,230 9,855 49,275 180
SARSim [SARSim_1, SARSim_2] 2017 Cls. 21,168 105,840
OpenSARShip [OpenSARship] 2017 Cls. 26,679 133,395
SAMPLE [SAMPLE] 2019 Cls. 5,380 26,900
ATRNet-STAR [ATRNet-STAR] 2025 Cls.&Det. 68,091 340,455 29,284 146,420 6,667
SARDet-100K [SARDet-100K] 2024 Det. 94,493 472,465 10,492 52,460 2,783
FAIR-CSAR [FAIR-CSAR] 2024 Det. 51,948 259,740 11,790 58,950 7,096
SARLANG-1M-Captions [SARLANG-1M] 2025 Cap. 9,191 31,968 3,939 13,682 3,902
SARVLM-1M(ours) 2025 Cap. 279,996 1,385,993 65,360 320,787 20,628
Figure 4: The framework of SARCap method.
3 Methodology
@@
This section introduce the paradigm of our framework, dataset construction approach and training strategy for SARCLIP and SARCap.
3.1 Problem Definition
@@
In this section, we investigate the paradigm of learning joint representations from SAR images and their corresponding textual descriptions. Specifically, we construct SARVLM-1M dataset $\mathcal{D}=\{(\mathbf{I}_{i},\mathbf{T}_{i})\}_{i=1}^{M}$ consisting of SAR images $\mathbf{I}_{i}\in\mathcal{R}^{H\times W}$ with the corresponding descriptions $\mathbf{T}_{i}\in\mathcal{T}$. As shown in Figure 3, our objective is to learn a pair of modality-specific encoders that project both SAR images and text into a shared semantic space. Specifically, we define a visual encoder $f_{v}:\mathcal{R}^{H\times W}\rightarrow\mathcal{R}^{d}$ that maps the input SAR image to a $d$-dimensional visual feature embedding $z_{v}^{i}=f_{v}({\mathbf{I}_{i}})$, and a textual encoder $f_{t}:\mathcal{R}^{\mathcal{T}}\rightarrow\mathcal{R}^{d}$ that maps the textual input to a corresponding textual embedding $z_{t}^{i}=f_{t}({\mathbf{T}_{i}})$. The goal is to align the embeddings $z_{v}^{i}$ and $z_{t}^{i}$ of matched image-text pairs in a common representation space, such that semantically similar inputs across modalities are embedded close to each other. This formulation enables the model to bridge the modality gap between SAR images and natural language, thereby supporting a wide range of downstream tasks such as cross-modal retrieval, target recognition, and zero-shot classification in SAR domains.
Figure 5: The word cloud of SARVLM-1M dataset.
3.2 SARVLM-1M Dataset Construction
@@
To address the aforementioned challenges, we build a large-scale vision-language dataset SARVLM-1M collected from existing classification, detection and caption datasets in SAR domain. As illustrated in Table 1, several large datasets have released in the past two year. For classification datasets, we employ 10 simple and complex description template to generate image captions. As for detection, we additionally design template to describe object absolute position and relative spatial position between targets in the images. The details for template design method are as follows:
@@
Based on the above templates design, we obtain caption for each SAR imagery. Then, A large language model is utilized to verify the fluency and grammatical correctness of text. Finally, we build our SARVLM-1M dataset contains 279,996 images and 1,385,993 captions in training set, and 20,628 image-text pairs in testing set. As shown in Figure 5, the dataset includes diverse target types such as ship, vehicle, aircraft, bridge and other land cover types.
3.3 Model Training
3.3.1 SARCLIP training
To further enhance the semantic understanding ability of the model for SAR image, we utilize two stages training strategy to train our SARCLIP. Firstly, we train vision large model on natural remote sensing dataset like RemoteCLIP which contains LoveDA [LoveDA], DOTA [DOTA], RSCID [RSCID] and others. By this strategy, our model is capable of better understanding and recognizing the unique characteristics of remote sensing imagery, particularly in identifying small scale and densely distributed objects. Secondly, we finetune this pretrained model on our SARVLM-1M dataset to transfer knowledge from natural to SAR domain.